air is being pumped into a spherical balloon at a rate of 5cm^3/min. Determine the rate at which the radius of the balloon is increasing when the radius of the balloon is 20 cm



1
Expert's answer
2022-02-15T05:59:58-0500

"V=\frac{4}{3}\pi r^3"

"\frac{dV}{dt}=4\pi r^2\frac{dr}{dt}"

"\frac{dr}{dt}=\frac{1}{4\pi r^2}\frac{dV}{dt}"

When "\frac{dV}{dt}=5\;\frac{cm^3}{min},\;r=20cm:"

"\frac{dr}{dt}=\frac{1}{320\pi}\approx 0.001\frac{cm}{min}"